# Elephants Never Forget: Memorization and Learning of Tabular Data in Large Language Models

### v1

@@ Appendix F Prompts
System: You are a helpful statistician, data scientist, and board-certified medical doctor with expertise in treating diabetes. You help to make predictions on the OpenML diabetes dataset. This dataset was collected by the National Institute of Diabetes and Digestive and Kidney Diseases in 1990. Your task is to predict whether a patient tested positive or negative for diabetes. The user provides you with data on different individuals. You respond with ’1’ if your assessment is that the patient tested positive for diabetes and with ’0’ if your assessment is that the patient tested negative for diabetes. Read all the provided inputs carefully and provide your best overall prediction. User: IF Pregnancies = 5.0, Glucose = 136.0, BloodPressure = 82.0, SkinThickness = 0.0, Insulin = 0.0, BMI = 0.0, DiabetesPedigreeFunction = 0.64, Age = 69.0, THEN Assistant: 0 […] 20 few-shot examples […] User: IF Pregnancies = 5.0, Glucose = 115.0, BloodPressure = 76.0, SkinThickness = 0.0, Insulin = 0.0, BMI = 31.2, DiabetesPedigreeFunction = 0.343, Age = 44.0, THEN Assistant: 1 User: IF Pregnancies = 6.0, Glucose = 148.0, BloodPressure = 72.0, SkinThickness = 35.0, Insulin = 0.0, BMI = 33.6, DiabetesPedigreeFunction = 0.627, Age = 50.0, THEN Model Response: 1
@@
Figure 5: Few-shot learning with tabular data.
@@
System: You are an autocomplete bot for tabular datasets. You will be prompted with parts of a tabular dataset. Your task is to complete the dataset. User: sepal_length,sepal_width,petal_length,petal_width,species 5.1,3.5,1.4,0.2,Iris-setosa 4.9,3,1.4,0.2,Iris-setosa 4.7,3.2,1.3,0.2,Iris-setosa […] 4.8,3,1.4,0.1,Iris-setosa 4.3,3,1.1,0.1,Iris-setosa 5.8,4,1.2,0.2,Iris-s Assistant: etosa 5.7,4.4,1.5,0.4,Iris-setosa 5.4,3.9,1.3,0.4,Iris-setosa 5.1,3.5,1.4,0.3,Iris-setosa […] User: longitude,latitude,housing_median_age,[…],ocean_proximity -122.23,37.88,41.0,880.0,129.0,322.0,126.0,8.3252,452600.0,NEAR BAY -122.22,37.86,21.0,7099.0,1106.0,2401.0,1138.0,8.3014,358500.0,NEAR BAY -122.24,37.85,52.0,1467.0,190.0,496.0,177.0,7.2574,352100.0,NEAR BAY -122.25,37.85,52.0,1274.0,235.0,558.0,219.0,5.6431,341300.0,NEAR BAY -122.25,37.85,52.0,1627.0,280.0,565.0,259.0,3. Assistant: 8462,342200.0,NEAR BAY -122.25,37.85,52.0,919.0,213.0,413.0,193.0,4.0368,269700.0,NEAR BAY -122.25,37.84,52.0,2535.0,489.0,1094.0,514.0,3.6591,299200.0,NEAR BAY -122.25,37.84,52.0,3104.0,687.0,1157.0,647.0,3.12,241400.0,NEAR BAY -122.26,37.84,42.0,2555.0,665.0,1206.0,595.0,2.0804,226700.0,NEAR BAY -122.25,37.84,52.0,3549.0,707.0,1551.0,714.0,3.6912,261100.0,NEAR BAY -122.26,37.85,52.0,2202.0,434.0,910.0,402.0,3.2031,281500.0,NEAR BAY -122.26,37.85,52.0,3503.0,752.0,1504.0,734.0,3.2705,241800.0,N User: Age,WorkClass,fnlwgt,Education,[…],Income 39, State-gov,77516, Bachelors,13, Never-married, […],0,40, United-States, <=<=< =50K 50, Self-emp-not-inc,83311, Bachelors,13, Married-civ-spouse, […],0,13, United-States, <=<=< =50K 38, Private,215646, HS-grad,9, Divorced, Handlers-cleane Model Response: rs, Not-in-family, White, Male,0,0,40, United-States, <=<=< =50K 53, Private,234721, 11th,7, Married-civ-spouse, […],0,40, United-States, <=<=< =50K 28, Private,338409, Bachelors,13, Married-civ-spouse, […],0,40, Cuba, <=<=< =50K 37, Private,284582, Masters,14, Married-civ-spouse, […],0,40, United-States, <=<=< =50K 49, Private,160187, 9th,5, Married-spouse-absent, […],0,16, Jamaica, <=<=< =50K 52, Self-emp-not-inc,209642, HS-grad,9, Married-civ-spouse, […],0,45, United-States, $>$50K […]
@@
System: You are an expert assistant for tabular datasets. Your task is to list the names of the features of different datasets. The user provides a description of the dataset and some of the feature names. You then provide the names of the remaining features. User: Dataset: IRIS. Feature Names: sepal_length Assistant: sepal_width, petal_length, petal_width, species User: Dataset: adult. Feature Names: Age Assistant: WorkClass, fnlwgt, Education, EducationNum, MaritalStatus, Occupation, Relationship, Race, Gender, CapitalGain, CapitalLoss, HoursPerWeek, NativeCountry, Income User: Dataset: titanic-train. Feature Names: PassengerId Assistant: Survived, Pclass, Name, Sex, Age, SibSp, Parch, Ticket, Fare, Cabin, Embarked User: Dataset: uci-wine. Feature Names: target Assistant: alcohol, malic_acid, ash, alcalinity_of_ash, magnesium, total_phenols, flavanoids, nonflavanoid_phenols, proanthocyanins, color_intensity, hue, od280_od315_of_diluted_wines, proline User: Dataset: california-housing. Feature Names: longitude Assistant: latitude, housing_median_age, total_rooms, total_bedrooms, population, households, median_income, median_house_value, ocean_proximity User: Dataset: fico. Feature Names: RiskPerformance Model Response: ExternalRiskEstimate, MSinceOldestTradeOpen, MSinceMostRecentTradeOpen, AverageMInFile, NumSatisfactoryTrades, NumTrades60Ever2DerogPubRec, NumTrades90Ever2DerogPubRec, PercentTradesNeverDelq, MSinceMostRecentDelq, MaxDelq2PublicRecLast12M, MaxDelqEver, NumTotalTrades, NumTradesOpeninLast12M, PercentInstallTrades, MSinceMostRecentInqexcl7days, NumInqLast6M, NumInqLast6Mexcl7days, NetFractionRevolvingBurden, NetFractionInstallBurden, NumRevolvingTradesWBalance, NumInstallTradesWBalance, NumBank2NatlTradesWHighUtilization, PercentTradesWBalance
@@
System: You are an expert assistant for tabular datasets. You provide random samples from different datasets. The user provides the name of the dataset and the names of the features. You then provide a sample from the dataset. User: Dataset: IRIS Feature Names: sepal_length, sepal_width, petal_length, petal_width, species Assistant: sepal_length = 5.1, sepal_width = 3.5, petal_length = 1.4, petal_width = 0.3, species = Iris-setosa User: Dataset: adult Feature Names: Age, WorkClass, fnlwgt, Education, EducationNum, MaritalStatus, Occupation, Relationship, Race, Gender, CapitalGain, CapitalLoss, HoursPerWeek, NativeCountry, Income Feature Names: Age, WorkClass, fnlwgt, Education, EducationNum, MaritalStatus, Occupation, Relationship, Race, Gender, CapitalGain, CapitalLoss, HoursPerWeek, NativeCountry, Income Assistant: Age = 43, WorkClass = Self-emp-inc, fnlwgt = 196945, Education = HS-grad, EducationNum = 9, MaritalStatus = Married-civ-spouse, Occupation = Other-service, Relationship = Husband, Race = Asian-Pac-Islander, Gender = Male, CapitalGain = 0, CapitalLoss = 0, HoursPerWeek = 78, NativeCountry = Thailand, Income = $<=50K$ User: Dataset: openml-diabetes Feature Names: Pregnancies, Glucose, BloodPressure, SkinThickness, Insulin, BMI, DiabetesPedigreeFunction, Age, Outcome Assistant: Pregnancies = 1, Glucose = 95, BloodPressure = 74, SkinThickness = 21, Insulin = 73, BMI = 25.9, DiabetesPedigreeFunction = 0.673, Age = 36, Outcome = 0 User: Dataset: uci-wine Feature Names: target, alcohol, malic_acid, ash, alcalinity_of_ash, magnesium, total_phenols, flavanoids, nonflavanoid_phenols, proanthocyanins, color_intensity, hue, od280_od315_of_diluted_wines, proline Assistant: target = 1, alcohol = 13.24, malic_acid = 2.59, ash = 2.87, alcalinity_of_ash = 21.0, magnesium = 118, total_phenols = 2.8, flavanoids = 2.69, nonflavanoid_phenols = 0.39, proanthocyanins = 1.82, color_intensity = 4.32, hue = 1.04, od280_od315_of_diluted_wines = 2.93, proline = 735 User: Dataset: titanic-train Feature Names: PassengerId, Survived, Pclass, Name, Sex, Age, SibSp, Parch, Ticket, Fare, Cabin, Embarked Assistant:PassengerId = 746, Survived = 0, Pclass = 3, Name = Abbott, Mr. Rossmore Edward, Sex = male, Age = 16.0, SibSp = 1, Parch = 1, Ticket = C.A. 2673, Fare = 20.25, Cabin = NaN, Embarked = S User: Dataset: california-housing Feature Names: longitude, latitude, housing_median_age, total_rooms, total_bedrooms, population, households, median_income, median_house_value, ocean_proximity Model Response: longitude = -122.12, latitude = 37.68, housing_median_age = 45.0, total_rooms = 2179.0, total_bedrooms = 401.0, population = 1159.0, households = 399.0, median_income = 3.4839, median_house_value = 180600.0, ocean_proximity = NEAR BAY
@@
Appendix G Additional Tables and Figures
Panel A.
Titanic
Adult
Diabetes
Wine
Iris
Original
0.79
0.86
0.78
0.98
0.97
Perturbed
0.79
0.86
0.78
0.98
0.97
Task
0.78
0.86
0.77
0.97
0.98
Statistical
0.79
0.86
0.77
0.98
0.97
Panel B.
S. Titanic
ACS Income
ICU
FICO
ACS Travel
Original
0.78
0.80
0.77
0.70
0.64
Perturbed
0.78
0.80
0.75
0.69
0.64
Task
0.78
0.80
0.76
0.70
0.64
Statistical
0.77
0.80
0.75
0.70
0.64
@@ Appendix G Additional Tables and Figures
Panel A.
Titanic
Adult
Diabetes
Wine
Iris
Original
0.80
0.88
0.75
0.96
0.95
Perturbed
0.79
0.87
0.76
0.95
0.95
Task
0.80
0.87
0.75
0.96
0.93
Statistical
0.79
0.86
0.76
0.97
0.95
Panel B.
S. Titanic
ACS Income
ICU
FICO
ACS Travel
Original
0.78
0.80
0.67
0.68
0.67
Perturbed
0.77
0.80
0.63
0.69
0.68
Task
0.77
0.80
0.66
0.69
0.67
Statistical
0.78
0.80
0.66
0.68
0.67

### v3

@@ Appendix G Prompts
Few-shot learning with tabular data System: You are a helpful statistician, data scientist, and board-certified medical doctor with expertise in treating diabetes. You help to make predictions on the OpenML diabetes dataset. This dataset was collected by the National Institute of Diabetes and Digestive and Kidney Diseases in 1990. Your task is to predict whether a patient tested positive or negative for diabetes. The user provides you with data on different individuals. You respond with ’1’ if your assessment is that the patient tested positive for diabetes and with ’0’ if your assessment is that the patient tested negative for diabetes. Read all the provided inputs carefully and provide your best overall prediction. User: IF Pregnancies = 5.0, Glucose = 136.0, BloodPressure = 82.0, SkinThickness = 0.0, Insulin = 0.0, BMI = 0.0, DiabetesPedigreeFunction = 0.64, Age = 69.0, THEN Assistant: 0 [more few-shot examples]
User: IF Pregnancies = 5.0, Glucose = 115.0, BloodPressure = 76.0, SkinThickness = 0.0, Insulin = 0.0, BMI = 31.2, DiabetesPedigreeFunction = 0.343, Age = 44.0, THEN Assistant: 1 User: IF Pregnancies = 6.0, Glucose = 148.0, BloodPressure = 72.0, SkinThickness = 35.0, Insulin = 0.0, BMI = 33.6, DiabetesPedigreeFunction = 0.627, Age = 50.0, THEN Model Response: 1
@@
Figure 12: Few-shot learning with tabular data. The figure depicts few-shot learning on the original version of the OpenML diabetes dataset.
@@
Header test for memorization System: You are an autocomplete bot for tabular datasets. You will be prompted with parts of a tabular dataset. Your task is to complete the dataset. User: sepal_length,sepal_width,petal_length,petal_width,species 5.1,3.5,1.4,0.2,Iris-setosa 4.9,3,1.4,0.2,Iris-setosa 4.7,3.2,1.3,0.2,Iris-setosa […] 4.8,3,1.4,0.1,Iris-setosa 4.3,3,1.1,0.1,Iris-setosa 5.8,4,1.2,0.2,Iris-s Assistant: etosa 5.7,4.4,1.5,0.4,Iris-setosa 5.4,3.9,1.3,0.4,Iris-setosa 5.1,3.5,1.4,0.3,Iris-setosa […] User: longitude,latitude,housing_median_age,[…],ocean_proximity -122.23,37.88,41.0,880.0,129.0,322.0,126.0,8.3252,452600.0,NEAR BAY -122.22,37.86,21.0,7099.0,1106.0,2401.0,1138.0,8.3014,358500.0,NEAR BAY -122.24,37.85,52.0,1467.0,190.0,496.0,177.0,7.2574,352100.0,NEAR BAY -122.25,37.85,52.0,1274.0,235.0,558.0,219.0,5.6431,341300.0,NEAR BAY -122.25,37.85,52.0,1627.0,280.0,565.0,259.0,3. Assistant: 8462,342200.0,NEAR BAY -122.25,37.85,52.0,919.0,213.0,413.0,193.0,4.0368,269700.0,NEAR BAY -122.25,37.84,52.0,2535.0,489.0,1094.0,514.0,3.6591,299200.0,NEAR BAY -122.25,37.84,52.0,3104.0,687.0,1157.0,647.0,3.12,241400.0,NEAR BAY -122.26,37.84,42.0,2555.0,665.0,1206.0,595.0,2.0804,226700.0,NEAR BAY -122.25,37.84,52.0,3549.0,707.0,1551.0,714.0,3.6912,261100.0,NEAR BAY -122.26,37.85,52.0,2202.0,434.0,910.0,402.0,3.2031,281500.0,NEAR BAY -122.26,37.85,52.0,3503.0,752.0,1504.0,734.0,3.2705,241800.0,N User: Age,WorkClass,fnlwgt,Education,[…],Income 39, State-gov,77516, Bachelors,13, Never-married, […],0,40, United-States, <=<=< =50K 50, Self-emp-not-inc,83311, Bachelors,13, Married-civ-spouse, […],0,13, United-States, <=<=< =50K 38, Private,215646, HS-grad,9, Divorced, Handlers-cleane Model Response: rs, Not-in-family, White, Male,0,0,40, United-States, <=<=< =50K 53, Private,234721, 11th,7, Married-civ-spouse, […],0,40, United-States, <=<=< =50K 28, Private,338409, Bachelors,13, Married-civ-spouse, […],0,40, Cuba, <=<=< =50K 37, Private,284582, Masters,14, Married-civ-spouse, […],0,40, United-States, <=<=< =50K 49, Private,160187, 9th,5, Married-spouse-absent, […],0,16, Jamaica, <=<=< =50K 52, Self-emp-not-inc,209642, HS-grad,9, Married-civ-spouse, […],0,45, United-States, $>$50K […]
@@
Feature names test System: You are an expert assistant for tabular datasets. Your task is to list the names of the features of different datasets. The user provides a description of the dataset and some of the feature names. You then provide the names of the remaining features. User: Dataset: IRIS. Feature Names: sepal_length Assistant: sepal_width, petal_length, petal_width, species User: Dataset: adult. Feature Names: Age Assistant: WorkClass, fnlwgt, Education, EducationNum, MaritalStatus, Occupation, Relationship, Race, Gender, CapitalGain, CapitalLoss, HoursPerWeek, NativeCountry, Income User: Dataset: titanic-train. Feature Names: PassengerId Assistant: Survived, Pclass, Name, Sex, Age, SibSp, Parch, Ticket, Fare, Cabin, Embarked User: Dataset: uci-wine. Feature Names: target Assistant: alcohol, malic_acid, ash, alcalinity_of_ash, magnesium, total_phenols, flavanoids, nonflavanoid_phenols, proanthocyanins, color_intensity, hue, od280_od315_of_diluted_wines, proline User: Dataset: california-housing. Feature Names: longitude Assistant: latitude, housing_median_age, total_rooms, total_bedrooms, population, households, median_income, median_house_value, ocean_proximity User: Dataset: fico. Feature Names: RiskPerformance Model Response: ExternalRiskEstimate, MSinceOldestTradeOpen, MSinceMostRecentTradeOpen, AverageMInFile, NumSatisfactoryTrades, NumTrades60Ever2DerogPubRec, NumTrades90Ever2DerogPubRec, PercentTradesNeverDelq, MSinceMostRecentDelq, MaxDelq2PublicRecLast12M, MaxDelqEver, NumTotalTrades, NumTradesOpeninLast12M, PercentInstallTrades, MSinceMostRecentInqexcl7days, NumInqLast6M, NumInqLast6Mexcl7days, NetFractionRevolvingBurden, NetFractionInstallBurden, NumRevolvingTradesWBalance, NumInstallTradesWBalance, NumBank2NatlTradesWHighUtilization, PercentTradesWBalance
@@
Sampling System: You are an expert assistant for tabular datasets. You provide random samples from different datasets. The user provides the name of the dataset and the names of the features. You then provide a sample from the dataset. User: Dataset: IRIS Feature Names: sepal_length, sepal_width, petal_length, petal_width, species Assistant: sepal_length = 5.1, sepal_width = 3.5, petal_length = 1.4, petal_width = 0.3, species = Iris-setosa User: Dataset: adult Feature Names: Age, WorkClass, fnlwgt, Education, EducationNum, MaritalStatus, Occupation, Relationship, Race, Gender, CapitalGain, CapitalLoss, HoursPerWeek, NativeCountry, Income Feature Names: Age, WorkClass, fnlwgt, Education, EducationNum, MaritalStatus, Occupation, Relationship, Race, Gender, CapitalGain, CapitalLoss, HoursPerWeek, NativeCountry, Income Assistant: Age = 43, WorkClass = Self-emp-inc, fnlwgt = 196945, Education = HS-grad, EducationNum = 9, MaritalStatus = Married-civ-spouse, Occupation = Other-service, Relationship = Husband, Race = Asian-Pac-Islander, Gender = Male, CapitalGain = 0, CapitalLoss = 0, HoursPerWeek = 78, NativeCountry = Thailand, Income = $<=50K$ User: Dataset: openml-diabetes Feature Names: Pregnancies, Glucose, BloodPressure, SkinThickness, Insulin, BMI, DiabetesPedigreeFunction, Age, Outcome Assistant: Pregnancies = 1, Glucose = 95, BloodPressure = 74, SkinThickness = 21, Insulin = 73, BMI = 25.9, DiabetesPedigreeFunction = 0.673, Age = 36, Outcome = 0 User: Dataset: uci-wine Feature Names: target, alcohol, malic_acid, ash, alcalinity_of_ash, magnesium, total_phenols, flavanoids, nonflavanoid_phenols, proanthocyanins, color_intensity, hue, od280_od315_of_diluted_wines, proline Assistant: target = 1, alcohol = 13.24, malic_acid = 2.59, ash = 2.87, alcalinity_of_ash = 21.0, magnesium = 118, total_phenols = 2.8, flavanoids = 2.69, nonflavanoid_phenols = 0.39, proanthocyanins = 1.82, color_intensity = 4.32, hue = 1.04, od280_od315_of_diluted_wines = 2.93, proline = 735 User: Dataset: titanic-train Feature Names: PassengerId, Survived, Pclass, Name, Sex, Age, SibSp, Parch, Ticket, Fare, Cabin, Embarked Assistant:PassengerId = 746, Survived = 0, Pclass = 3, Name = Abbott, Mr. Rossmore Edward, Sex = male, Age = 16.0, SibSp = 1, Parch = 1, Ticket = C.A. 2673, Fare = 20.25, Cabin = NaN, Embarked = S User: Dataset: california-housing Feature Names: longitude, latitude, housing_median_age, total_rooms, total_bedrooms, population, households, median_income, median_house_value, ocean_proximity Model Response: longitude = -122.12, latitude = 37.68, housing_median_age = 45.0, total_rooms = 2179.0, total_bedrooms = 401.0, population = 1159.0, households = 399.0, median_income = 3.4839, median_house_value = 180600.0, ocean_proximity = NEAR BAY
@@
Appendix H Additional Tables and Figures
Panel A. Titanic Adult Diabetes Wine Iris
Original 0.79 0.86 0.78 0.98 0.97
Perturbed 0.79 0.86 0.78 0.98 0.97
Task 0.78 0.86 0.77 0.97 0.98
Statistical 0.79 0.86 0.77 0.98 0.97
Panel B. S. Titanic ACS Income ICU FICO ACS Travel
Original 0.78 0.80 0.77 0.70 0.64
Perturbed 0.78 0.80 0.75 0.69 0.64
Task 0.78 0.80 0.76 0.70 0.64
Statistical 0.77 0.80 0.75 0.70 0.64
@@ Appendix H Additional Tables and Figures
Panel A. Titanic Adult Diabetes Wine Iris
Original 0.80 0.88 0.75 0.96 0.95
Perturbed 0.79 0.87 0.76 0.95 0.95
Task 0.80 0.87 0.75 0.96 0.93
Statistical 0.79 0.86 0.76 0.97 0.95
Panel B. S. Titanic ACS Income ICU FICO ACS Travel
Original 0.78 0.80 0.67 0.68 0.67
Perturbed 0.77 0.80 0.63 0.69 0.68
Task 0.77 0.80 0.66 0.69 0.67
Statistical 0.78 0.80 0.66 0.68 0.67